a page with length 3 inches more than its width, has 80 square inches of printed area. the margins at the top and bottom are each 1 inch wide. Find the dimension of the page
Answer by ikleyn(52781)

.
Let w be the page width, in inches.


Then its length is (w+3) inches.


The printed area has dimensions in 2 = 1 + 1 inches less. Therefore, the equation for the printed area is


(w-2)*(w+1) = 80.


w^2 -w - 82 = 0.


Solve it using the quadratic formula.

------------------

To see many other similar solved problems, look into the lesson
    - Problems on the area and the dimensions of a rectangle surrounded by a strip
    - Cynthia Besch wants to buy a rug for a room
in this site.